# Dynamics of the sine-Gordon equation on tadpole graphs

### v1

@@ Abstract
This work studies the dynamics of solutions to the sine-Gordon equation posed on a tadpole graph $\mathcal{G}$ and endowed with boundary conditions at the vertex of $\delta$-type. The latter generalize conditions of Neumann-Kirchhoff type. The purpose of this analysis is to establish an instability result for a certain family of stationary solutions known as single-lobe kink state profiles, which consist of a periodic, symmetric, concave stationary solution in the finite (periodic) lasso of the tadpole, coupled with a decaying kink at the infinite edge of the graph. It is proved that such stationary profile solutions are linearly (and nonlinearly) unstable under the flow of the sine-Gordon model on the graph. The extension theory of symmetric operators, Sturm-Liouville oscillation results and analytic perturbation theory of operators are fundamental ingredients in the stability analysis. The local well-posedness of the sine-Gordon model in an appropriate energy space is also established. The theory developed in this investigation constitutes the first stability result of stationary solutions to the sine-Gordon equation on a tadpole graph.
@@ 2 Instituto de Investigaciones en Matemáticas Aplicadas y en Sistemas,
Key words. sine-Gordon model, metric graphs, kink and anti-kink solutions, $\delta$-type interaction, analytic perturbation theory, extension theory, instability.
@@ 1 Introduction
Nonlinear evolution models on quantum graphs emerge as useful tools for studying wave propagation, particularly in quasi-one-dimensional systems such as meso- or nanoscale structures that resemble a thin neighborhood around a graph. These models offer a practical way to explore various physical phenomena in both theoretical and experimental contexts. The flexibility in configuring the graph’s geometry enables the generation of diverse dynamics. Examples of bifurcated systems include Josephson junction structures, networks of planar optical waveguides and fibers, branched structures like DNA, blood pressure waves in large arteries, nerve impulses in complex neuron arrays, conducting polymers, and Bose-Einstein condensation (see, e.g., [20, 23, 24, 26, 27, 34, 36, 41, 48] and references therein).
Recently, nonlinear models on metric graphs, such as the nonlinear Schrödinger equation, the sine-Gordon model, and the Korteweg-de Vries equation have garnered considerable attention (for an abridged list of related works the reader is referred to [1, 12, 11, 8, 9, 14, 15, 16, 2, 27, 41]). Mathematically, an evolution model on a graph is equivalent to a system of partial differential equations (PDEs) defined along the edges (viewed as intervals), with the coupling determined solely by the boundary conditions at the vertices—this is referred to as the “topology of the graph”. These boundary conditions govern the network’s dynamics through group interactions. In particular, the study of soliton transport has been particularly prominent in this field, as solitons and other nonlinear waves in branched systems provide valuable insights into the behavior of these models.
Addressing these challenges is difficult due to the complexity of both the equations of motion and the graph topology. A particular issue complicating the analysis is the behavior at the vertices, where a soliton profile traveling along one edge may experience complex phenomena like reflection and radiation emission. This makes it hard to track how energy propagates through the network. Consequently, studying soliton propagation in networks presents significant obstacles. The mechanisms governing the existence and stability (orbital or spectral) of soliton profiles remain uncertain for many types of graphs and models. This project intends to contribute with the exploration of these topics in the particular case of the sine-Gordon equation,
@@
We recall that a metric graph $\mathcal{G}$ is a structure represented by a finite number of vertices $V=\{\nu_{i}\}$ and a set of adjacent edges at the vertices $E=\{e_{j}\}$ (for further details, see [20]). Each edge $e_{j}$ can be identified with a finite or infinite interval of the real line, $I_{e}$. Thus, the edges of $\mathcal{G}$ are not merely abstract relations between vertices but can be viewed as physical “wires” or “networks” connecting them. The notation $e\in E$ will be used to indicate that $e$ is an edge of $\mathcal{G}$. This identification introduces the coordinate $x_{e}$ along the edge $e$. A quantum graph is a metric graph $\mathcal{G}$ equipped with a differential (or pseudo-differential) operator acting on functions on the graph.
One of the main graphs of interest to us here is the so-called tadpole graph (also known as lasso graph [26]), a graph composed of a circle and a infinite half-line attached to a common vertex $\nu=L$. If the circle is identified with the interval $[-L,L]$ and the half-line with $[L,\infty)$, we obtain a particular metric graph structure that we will denote again by $\mathcal{G}$, represented by $E=\{[-L,L],[L,\infty)\}$ and $V=\{L\}$ (see Figure 1). We use the subscripts $1$ and $2$ to refer to the edges $e_{1}=[-L,L]$ and $e_{2}=[L,\infty)$.
@@
A wave function U defined on the tadpole graph $\mathcal{G}$ will be understood as an pair of functions $\textbf{U}=(u_{1},u_{2})$, where $u_{1}$ is defined on $E_{1}=[-L,L]$ and $u_{2}$ on $E_{2}=[L,\infty)$. Thus, we define the action of the Laplacian operator $\Delta$ on the tadpole graph $\mathcal{G}$ by
$-\Delta:(u_{e})_{e\in E}\to(-u^{\prime\prime}_{e})_{e\in E}.$ (1.2)
Thus, from (1.1), for $\mathbf{U}(x_{e},t)=(u_{1}(x_{e},t),u_{2}(x_{e},t))$ and the nonlinearity $\sin\mathbf{U}$ acting componentwise, i.e., $\sin\mathbb{U}(x_{e},t)=(\sin u_{1}(x_{e},t),\sin u_{2}(x_{e},t))$, we obtain that for $\textbf{c}^{2}=(c_{1}^{2},c_{2}^{2})$ the model in (1.1) can be written in the following vectorial form on a tadpole graph $\mathcal{G}$,
@@
Posing the sine-Gordon equation on a metric graph comes out naturally from practical applications. For example, since the phase-difference in a long (infinite) Josephson junction obeys the classical sine-Gordon model on the line (cf. [18, 32, 50]), the coupling of two or more Josephson junctions forming a network can be effectively modeled by the vectorial sine-Gordon model in (1.1) on a graph. The sine-Gordon equation was first conceived on a $\mathcal{Y}$-shaped Josephson junction by Nakajima et al. [40, 39] as a prototype for logic circuits. The authors consider three long (semi-infinite) Josephson junctions coupled at one single common vertex, a structure known as a tricrystal junction. There exists two main types of $\mathcal{Y}$-junctions. A $\mathcal{Y}$-junction of the first type (or type I) consists of one incoming (or parent) edge, $E_{1}=(-\infty,0)$, meeting at one single vertex at the origin, $\nu=0$, with other two outgoing (children) edges, $E_{j}=(0,\infty)$, $j=2,3$. The second type (or $\mathcal{Y}$-junction of type II) resembles more a starred structure and consists of three identical edges of the form $E_{j}=(0,\infty)$, $1\leqq j\leqq 3$. Junctions of type I are more common in unidirectional fluid flow models (see, for example, [21]), whereas graphs of type I or II are indistinctively used to describe Josephson tricrystal junctions; see, for instance, [29, 49] (type I), or [50, 33, 46] (type II). Recently, Angulo and Plaza [14, 15, 16] studied the sine-Gordon model on $\mathcal{Y}$-junctions of type I and II by obtaining an comprehensive study of the local well-posedness for the Cauchy problem and the linear (and nonlinear) instability results of static, kink and kink/anti-kink soliton profile solutions. The established theory considers boundary conditions at the graph-vertex of either $\delta$-interaction type (expressing continuity of the wave functions plus a Kirchhoff-type rule expressing conservation of the current low, the Kirchhoff law for electric currents) or $\delta^{\prime}$-interaction type (expressing continuity of the derivatives of the wave functions plus a Kirchhoff-type rule for the self-induced magnetic flux).
@@
[…]
@@
for all $j=1,2$, and $x\in E_{j}$, $t>0$, where each of the profile functions $\phi_{j}$ satisfies the equation
$-c_{j}^{2}\phi^{\prime\prime}_{j}+\sin\phi_{j}=0,$ (1.6)
on each edge $E_{j}$ and for all $j$, as well as the boundary conditions at the vertex $\nu=0$ being the so-called $\delta$-coupling or generalized Neumann-Kirchhoff boundary conditions, namely, $(\phi_{1},\phi_{2})\in D_{Z}$
@@
The parameter $Z\in\mathbb{R}$ is a coupling constant between the disconnected loop and the half-line. The values of $f$, $g$ and their derivatives at the vertex (and at $-L$) are understood as the appropriate one-sided limits (observe that, as $(f,g)\in H^{2}(-L,L)\times H^{2}(L,\infty)$, $f^{\prime}(\pm L)$ and $g^{\prime}(L+)$ are well defined). We note that $(-\Delta,D_{Z})_{Z\in\mathbb{R}}$ represents a one-parameter family of self-adjoint operators on a tadpole graph (see Theorem A.4 in the Appendix).
The delicate point about the existence of stationary solutions solving (1.6) is given by the component $\phi_{1}$ on $[-L,L]$. The profile $\phi_{2}$ on $[L,\infty)$ is given by the well-known kink-soliton profile on the full real line modulo translation (see [25, 47]),
@@
where we are interested for the case $c_{2}>0$ ($\lim_{x\to\infty}\phi_{2}(x)=0$) and the shift parameter $a\in\mathbb{R}$ will be determined by the condition $(\phi_{1},\phi_{2})\in D_{Z}$ (see Figure 2).
@@
Among all stationary solutions for (1.6) (specifically on the lasso, where we can see the existence of infinite profiles with many lobes), we are particularly interested in the so-called single-lobe kink states, an example of which is shown on Figure 3. More precisely, we have the following definition.
Figure 3: An example of the graph a single-lobe kink state profile for the sine-Gordon model according to Definition 1.1.
@@
The stationary profile solution $\Theta=(\phi_{1},\phi_{2})\in D_{Z}$ is said to be a single-lobe kink state for (1.6) if every component is positive on each edge of the tadpole graph $\mathcal{G}$, the maximum of $\phi_{1}$ is achieved at a single internal point, symmetric on $[-L,L]$ and monotonically decreasing on $[0,L]$. Moreover, $\phi_{2}$ is strictly decreasing on $[L,\infty)$.
@@
For the convenience of the reader we give a brief description of our main results. In Propositions 4.2 and 4.4 below, we establish the profiles of a two-families of single-lobe states for (1.6) on $[-L,L]$ depending of the signal of the kink-shift, $a<0$ or $a>0$ (but no necessarily being the first component for a single-lobe kink state $(\phi_{1},\phi_{2})\in D_{Z}$). In Proposition 4.2 ($a<0$), this family of solutions $\phi_{1}\equiv\phi_{1,k}$ for equation in (1.6) are a piece of subluminal periodic type solution for the sine-Gordon with a single-lobe profile on $[-L,L]$ given by,
@@
where $\operatorname{sn}(\cdot;k)$ is the Jacobian elliptic function of snoidal-type and the elliptic modulus $k$ satisfying $K(k)>\frac{L}{c_{1}}$. More precisely, for $\frac{L}{c_{1}}\leqq\frac{\pi}{2}$, we have $k\in(0,1)$, and for $\frac{L}{c_{1}}>\frac{\pi}{2}$, we need to have $k\in(k_{0},1)$, with $K(k_{0})=\frac{L}{c_{1}}$, by properties of the strictly increasing Legendre’s complete elliptic integral of the first type $K$ (see, e.g., Byrd and Friedman [22]). The profile $\phi_{1,k}$ satisfies $\phi^{\prime\prime}_{1,k}(x)<0$ for all $x\in[-L,L]$ (see Figure 4). Moreover, In Proposition 4.7 below we give a complete restrictions on the parameters $L,c_{1},c_{2},k$ and $Z$ for which there are (or there are not) single-lobe kink states $(\phi_{1,k},\phi_{2,a(k)})\in D_{Z}$.
@@
The second goal of the present analysis is associated with the dynamics of the sine-Gordon flow around the kink profiles obtained in Proposition 4.7; more specifically, we are interested on its stability properties. The stability of these static configurations is an important property from both the mathematical and the physical points of view. Stability can predict whether a particular state can be observed in experiments or not. Unstable configurations are rapidly dominated by dispersion, drift, or by other interactions depending on the dynamics, and they are practically undetectable in applications.
In the forthcoming stability analysis, the family of linearized operators around the single-lobe kink states $(\phi_{1},\phi_{2})\equiv(\phi_{1,k},\phi_{2,a(k)})\in D_{Z}$ determined by Proposition 4.7 plays a fundamental role. These operators are characterized by the following formal self-adjoint diagonal matrix operator,
@@
where $\delta_{j,k}$ denotes the Kronecker symbol, and defined on domains with $\delta$-type interaction at the vertex $\nu=0$, $D(\mathcal{L}_{Z})=D_{Z}$ in (1.7). It is to be observed that the particular family determined by Proposition 4.7 of single-lobe kink states under consideration is such that $(\phi_{1,k},\phi_{2,a(k)})\in D(\mathcal{L}_{Z})$ for every $k$ admissible. An interesting characteristic of the spectrum structure associated with operators in (1.10) on a tadpole graph is that they have a non-trivial Morse index (larger than or equal to 1, in general) which makes the stability study not so immediate. Here we use a novel linear instability criterion for stationary solutions of evolution models on metric graphs (see Theorem 3.2 below) developed by Angulo and Cavalcante [9] (see also [14, 15, 16]) and adapted to the case of the sine-Gordon model on tadpole graphs, as well as a splitting eigenvalue method introduced by Angulo [7] (see Lemma 5.1 below and Angulo [6, 7]). This method is applied to the operator $\mathcal{L}_{Z}\equiv\operatorname{diag}(\mathcal{L}_{1},\mathcal{L}_{a})$ on $D_{Z}$. More precisely, we reduce the eigenvalue problem associated to $\mathcal{L}_{Z}$ into two classes of eigenvalue problems, one for $\mathcal{L}_{1}$ with periodic boundary conditions on $[-L,L]$ and the second one for $\mathcal{L}_{a}$ with $\delta$-type boundary condition on the half-line $[L,\infty)$. Hence, under the restrictions on $L,c_{1},c_{2},Z$ and $k$ given in Theorems 5.10, 5.13 and 5.14, we show that $\ker(\mathcal{L}_{Z})=\{\mathbb{0}\}$ and that the Morse index of $\mathcal{L}_{Z}$ is equal to one. Therefore, as a consequence of the linear instability criterion in Theorem 3.2 we deduce that these single-lobe kink states are linearly unstable by the sine-Gordon flow on a tadpole graph (see Theorem 5.14 below).
@@
1.1 Main results
Let us summarize the main contributions of this paper:
    First, we prove that the Cauchy problem associated to (1.4) on a tadpole graph $\mathcal{G}$ is well-posed in the energy space $\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G})$ with $\mathcal{E}(\mathcal{G})$ defined in (2.1) (see Section 2). This is the content of Theorem 2.1. Even though the well-posedness result is not part of the spectral stability analysis, it is fundamental to reach a nonlinear conclusion; see Theorem 5.15 below.
    In Section 3 we establish a general instability criterion for static solutions for the sine-Gordon model (1.4) on a tadpole graph. The reader can find this result in Theorem 3.2 below. It essentially provides sufficient conditions on the flow of the semigroup generated by the linearization of (1.4) around a priori stationary solutions, for the existence of a pair of positive/negative real eigenvalues of the linearized operator based on its Morse and nullity indexes.
    We provide a complete characterization of the single-lobe kink state profiles $\Theta=(\phi_{1},\phi_{2})$ for the sine-Gordon model on a tadpole graph defined in Definition 1.1 in terms of the parameters $L,c_{1},c_{2}$ and $Z\in(-\infty,\frac{2}{\pi c_{2}})$, in the specific case of the shift-parameter $a<0$ in the kink-profile formula in (1.8). It is shown that $\Theta$ belongs to the domain space $D_{Z}$ in (1.7) only for specific relations between $L,c_{1},c_{2}$, $Z$, and the elliptic modulus $k$ defining profile $\phi_{1,k}$ in (1.9). This implies, in turn, that there does not exist a single-lobe kink state solution with a first-component being of the form in (1.9) (and hence, not satisfying the $\delta$-boundary condition in $D_{Z}$), for a large set of parameters $L,c_{1},c_{2},Z,k$ (see Proposition 4.7).
    We determine (see Section 5) the specific spectral structure for the family of self-adjoint operators $(\mathcal{L}_{Z},D_{Z})$ with $\mathcal{L}_{Z}$ defined in (1.10) and $(\phi_{1},\phi_{2})=(\phi_{1,k},\phi_{2,a(k)})\in D_{Z}$ determined by Proposition 4.7. In particular, we establish a splitting eigenvalue method in Lemma 5.1 which is an essential ingredient to prove that the Morse index of $\mathcal{L}_{Z}$ is exactly one and with a trivial kernel (see Theorem 5.14).
    In Theorem 5.15 we establish that the single-lobe kink state profiles $(\phi_{1,k},\phi_{2,a(k)})\in D_{Z}$ determined by Proposition 4.7, are spectrally and nonlinearly unstable for the sine-Gordon model (1.4) on a tadpole graph. As we mentioned above, the fact that the mapping data-solution associated to the sine-Gordon model on a tadpole graph is of class $C^{2}$ on the energy space $\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G})$ is crucial. We also obtained that the degenerate single-lobe kink state $(\pi,\phi_{2,0})\in D_{Z}$ with $Z=\frac{2}{\pi c_{2}}$, $\phi_{2,0}(x)=4\arctan\big{[}e^{-\frac{1}{c_{2}}(x-L)}\big{]}$, $x\in[L,\infty)$, is also nonlinearly unstable.
    For completeness and convenience of the reader, in Appendix A we formulate some tools from the extension theory of symmetric operators by Krein and von Neumann which are suitable for our needs. In particular, we establish a Perron-Frobenius result which is used in the accurate Morse index estimate $n(\mathcal{L}_{Z})\leqq 1$ for $\mathcal{L}_{Z}$ in (1.10) and $(\phi_{1},\phi_{2})$ being a priori single-lobe kink state. Moreover, for completeness in the exposition we also show the general estimate $n(\mathcal{L}_{Z})\leqq 2$ using extension theory.
@@ 1.1 Main results
[…]
@@
[…]

### v2

@@ Abstract
This work studies the dynamics of solutions to the sine-Gordon equation posed on a tadpole graph, namely, a graph consisting of a circle with a half-line attached at a single vertex, and endowed with boundary conditions at the vertex of $\delta$-type. The latter generalize conditions of Neumann-Kirchhoff type. The purpose of this analysis is to establish the existence and instability of stationary solutions which we have called single-lobe kink state profiles, which consist of a symmetric stationary solution of subluminal-type profile in the finite ring of the tadpole, coupled with a decaying kink-profile at the infinite edge of the graph. It is proved that such stationary profile solutions are linearly (and nonlinearly) unstable under specific restriction and by the flow of the sine-Gordon model on the graph. The extension theory of symmetric operators, Sturm-Liouville oscillation results and a splitting eigenvalue method are fundamental ingredients in the stability analysis. The local well-posedness of the sine-Gordon model in an appropriate energy space is also established. The theory developed in this investigation constitutes the first steps by understading the dynamics to the sine-Gordon equation on a tadpole graph.
@@ 2 Instituto de Investigaciones en Matemáticas Aplicadas y en Sistemas,
Key words. sine-Gordon model, tadpole graph, single-lobe kink solutions, $\delta$-type interaction, extension theory of symmetric operators, Sturm-Liouville theory, instability.
@@ 1 Introduction
Recently, nonlinear models posed on metric graphs, such as the nonlinear Schrödinger equation, the sine-Gordon model, and the Korteweg-de Vries equation, have garnered considerable attention (for an abridged list of related works the reader is referred to [1, 11, 10, 7, 8, 13, 14, 15, 2, 27, 28, 25, 26, 33]). Real systems can exhibit strong inhomogeneities due to varying nonlinear coefficients across different regions of the spatial domain or due to the specific geometry of the domain itself. Addressing these challenges is difficult due to the complexity of both the equations of motion and the graph topology. A particular issue complicating the analysis is the behavior at the vertices, where a soliton profile traveling along one edge may experience complex phenomena like reflection and radiation emission. This makes it hard to track how energy propagates through the network. Consequently, studying soliton propagation in networks presents significant obstacles. The mechanisms governing the existence and stability (orbital or spectral) of soliton profiles remain uncertain for many types of graphs and models. This project intends to contribute with the exploration of these topics in the particular case of the sine-Gordon equation,
@@
To explore key features of the sine-Gordon model on metric graphs we therefore choose a relatively simple metric graph domain known as the tadpole graph (see Figure 1), a study that has not been addressed in the literature, as far as we know.
We recall that a metric graph $\mathcal{G}$ is a structure represented by a finite number of vertices $V=\{\nu_{i}\}$ and a set of adjacent edges at the vertices $E=\{e_{j}\}$ (for further details, see [18]). Each edge $e_{j}$ can be identified with a finite or infinite interval of the real line, $I_{e}$. The notation $e\in E$ will be used to indicate that $e$ is an edge of $\mathcal{G}$. This identification introduces the coordinate $x_{e}$ along the edge $e$. A quantum graph is a metric graph $\mathcal{G}$ equipped with a differential (or pseudo-differential) operator acting on functions on the graph. A tadpole graph is a graph composed of a circle and a infinite half-line attached to a common vertex $\nu=L$. If the circle is identified with the interval [L,L][-L,L] and the half-line with $L,\infty)$, we obtain a particular metric graph structure that we will denote again by $\mathcal{G}$, represented by the edge-set $E=\{[-L,L],[L,\infty)\}$ and the unique vertex $V=\{L\}$ (see Figure 1). We use the subscripts $1$ and $2$ to refer to the edges $e_{1}=[-L,L]$ and $e_{2}=[L,\infty)$.
@@
A wave function U defined on the tadpole graph $\mathcal{G}$ will be understood as an pair of functions $\textbf{U}=(u_{1},u_{2})$, where $u_{1}$ is defined on $e_{1}$ and $u_{2}$ on $e_{2}$. Thus, we define the action of the Laplacian operator $\Delta$ on the tadpole graph $\mathcal{G}$ by
$\Delta\textbf{U}=(-u^{\prime\prime}_{1},-u^{\prime\prime}_{2}).$ (1.2)
Thus, from (1.1), for $\mathbf{U}(t)=(u_{1}(\cdot,t),u_{2}(\cdot,t))$ and the nonlinearity $\sin\mathbf{U}$ acting componentwise, i.e., $\sin\mathbb{U}(t)=(\sin u_{1}(\cdot,t),\sin u_{2}(\cdot,t))$, we obtain that for $\textbf{c}^{2}=(c_{1}^{2},c_{2}^{2})$ the model in (1.1) can be written in the following vectorial form on a tadpole graph $\mathcal{G}$,
@@
Posing the sine-Gordon equation on a metric graph comes out naturally from practical applications. For example, since the phase-difference in a long (infinite) Josephson junction obeys the classical sine-Gordon model on the line (cf. [16, 24, 39]), the coupling of three or more Josephson junctions forming a network can be effectively modeled by the vectorial sine-Gordon model in (1.1) on a graph. The sine-Gordon equation was first conceived on a $\mathcal{Y}$-shaped Josephson junction by Nakajima et al. [32, 31] as a prototype for logic circuits. Recently, Angulo and Plaza [13, 14, 15] studied the sine-Gordon model on $\mathcal{Y}$-junctions by obtaining an comprehensive study of the linear (and nonlinear) instability of static, kink and kink/anti-kink soliton profile solutions.
@@
$\begin{cases}\partial_{t}u_{j}=v_{j}\\ \partial_{t}v_{j}=c_{j}^{2}\partial_{x}^{2}u_{j}-\sin u_{j},\end{cases}\qquad x\in e_{j},\;\,t>0,\;\,1\leqq j\leqq 2$ (1.4)
@@
for all $j=1,2$, and $x\in e_{j}$, $t>0$, where each of the profile functions $\phi_{j}$ satisfies the classical kink-equation
$c_{j}^{2}\phi^{\prime\prime}_{j}+\sin\phi_{j}=0,$ (1.6)
on each edge $e_{j}$ and for all $j$, as well as the boundary conditions at the vertex $\nu=L$ being the so-called $\delta$-coupling or generalized Neumann-Kirchhoff boundary conditions, namely, $(\phi_{1},\phi_{2})\in D_{Z}$
@@
The parameter $Z\in\mathbb{R}$ is a coupling constant between the disconnected loop and the half-line. The values of $f$, $g$ and their derivatives at the vertex (and at L-L) are understood as the appropriate one-sided limits (observe that, as $(f,g)\in H^{2}(-L,L)\times H^{2}(L,\infty)$, $f^{\prime}(\pm L)$ and $g^{\prime}(L+)$ are well defined). We note that $(-\Delta,D_{Z})_{Z\in\mathbb{R}}$ represents a one-parameter family of self-adjoint operators on a tadpole graph via the extension theory of symmetric operators by Krein and von Neumann (see Theorem A.6 in [6]).
The delicate point about the existence of stationary solutions solving (1.6) is given by the component $\phi_{1}$ on [L,L][-L,L]. The profile $\phi_{2}$ on $L,\infty)$ is given by the well-known kink-soliton profile on the full real line modulo translation (see [20, 38]),
@@
where we are interested for the case $c_{2}>0$ ($\lim_{x\to+\infty}\phi_{2}(x)=0$) and the shift parameter $a\in\mathbb{R}$ will be determined by the condition $(\phi_{1},\phi_{2})\in D_{Z}$ (see Figure 2).
@@
Among all stationary solutions for (1.6), specifically on [L,L][-L,L], we are particularly interested in the so-called single-lobe kink states, an example of which is shown on Figure 3. More precisely, we have the following definition.
Figure 3: An example of a single-lobe kink state profile for the sine-Gordon model according to Definition 1.1.
@@
The stationary profile solution $\Theta=(\phi_{1},\phi_{2})\in D_{Z}$ is said to be a single-lobe kink state for (1.6) if every component is positive on each edge of the tadpole graph $\mathcal{G}$, the maximum of $\phi_{1}$ is achieved at a single internal point, symmetric on [L,L][-L,L] and monotonically decreasing on $[0,L]$. Moreover, $\phi_{2}$ is strictly decreasing on $L,\infty)$.
In Figure 6 (see formula (4.7)), we show other profile type for $\phi_{1}$ solution of (1.6) on [L,L][-L,L].
@@
For the convenience of the reader we give a brief description of our main results. Initially, in Propositions 4.2 and 4.4 below, we establish the profiles of a two-families of single-lobe states for (1.6) on [L,L][-L,L] depending on an a priori sign of the kink-shift, namely, $a<0$ or $a>0$ (we note that no necessarily all these profiles will be the first component for a single-lobe kink state $(\phi_{1},\phi_{2})\in D_{Z}$, such as is established in Propositions 4.7 and 4.11 below). In Proposition 4.2 (for a priori $a<0$), the following family of solutions $\phi_{1,k}$ for equation in (1.6) will be constructed based on the subluminal periodic type solution for the sine-Gordon (see Figure 5) with a single-lobe profile on [L,L][-L,L] given by,
@@
where $\operatorname{sn}(\cdot;k)$ is the Jacobian elliptic function of snoidal-type and the elliptic modulus $k$ satisfying $K(k)>\frac{L}{c_{1}}$, where $K$ is the Legendre’s complete elliptic integral of the first type (see (4.6) and Byrd and Friedman [19]). More precisely, for $\frac{L}{c_{1}}\leqq\frac{\pi}{2}$, we have the curve of solutions $k\in(0,1)\to\phi_{1,k}$, and for $\frac{L}{c_{1}}>\frac{\pi}{2}$, the curve $k\in(k_{0},1)\to\phi_{1,k}$, with $K(k_{0})=\frac{L}{c_{1}}$. The profile $\phi_{1,k}$ satisfies $\phi^{\prime\prime}_{1,k}(x)<0$ for all $x\in[-L,L]$ (see Figure 4). In this way, in Proposition 4.7 below we give a complete restrictions on the parameters $L,c_{1},c_{2},k$ and $Z$ for which there are (or there are not) single-lobe kink states $(\phi_{1,k},\phi_{2,a(k)})\in D_{Z}$ with a first-component being $\phi_{1,k}$ in (1.9). Similarly, for $a>0$, in Proposition 4.4 we establish also a family of solutions $\phi_{1,k}$ for equation in (1.6) (see (4.9)) but they are not strictly concave (see Figure 9). in Proposition 4.11, we also give sufficient conditions on $L,c_{1},c_{2},k$ and $Z\leqq 0$ for the subluminal profile in (4.9) to be the first-component of a single-lobe kink state. In Proposition 4.6 , we establish the existence of degenerated single-lobe kink state $(\pi,\phi_{2,0})\in D_{Z}$ if and only if $Z=\frac{2}{\pi c_{2}}$ (see Figure 10).
@@
The second goal of the present analysis is associated with the dynamics of the sine-Gordon flow around the single-lobe kink states obtained in Propositions 4.6, 4.7 and 4.11; more specifically, we are interested on its stability properties. The stability of these static configurations is an important property from both the mathematical and the physical points of view. Stability can predict whether a particular state can be observed in experiments or not. Unstable configurations are rapidly dominated by dispersion, drift, or by other interactions depending on the dynamics, and they are practically undetectable in applications.
In the forthcoming stability analysis, for instance, by considering the single-lobe kink states $(\phi_{1},\phi_{2})\equiv(\phi_{1,k},\phi_{2,a(k)})\in D_{Z}$ determined by Proposition 4.7 (with $\phi_{1,k}$ in (1.9)), the following family of linearized operators around of $(\phi_{1},\phi_{2})$ plays a fundamental role,
@@
where $\delta_{j,k}$ denotes the Kronecker symbol, and defined on domains with $\delta$-type interaction at the vertex $\nu=L$, $D(\mathcal{L}_{Z})=D_{Z}$ in (1.7). An interesting characteristic of the spectrum structure associated with operators in (1.10) on a tadpole graph is that they have a non-trivial Morse index (larger than or equal to 1, in general) which makes the stability study not so immediate. Here we use a novel linear instability criterion for stationary solutions of evolution models on metric graphs (see Theorem 3.2 below) developed by Angulo and Cavalcante [8] and adapted to the case of the sine-Gordon model on tadpole graphs, as well as a splitting eigenvalue method introduced by Angulo [6] (see also Lemma 5.1 below and Angulo [5]). This method is applied to the operator $\mathcal{L}_{Z}\equiv\operatorname{diag}(\mathcal{L}_{1},\mathcal{L}_{a})$ on $D_{Z}$. More precisely, we reduce the eigenvalue problem associated to $\mathcal{L}_{Z}$ in (1.10) into two classes of eigenvalue problems, one for $\mathcal{L}_{1}$ with periodic boundary conditions on [L,L][-L,L] and the second one for $\mathcal{L}_{a}$ with $\delta$-type boundary condition on the half-line $L,\infty)$. Hence, under the restrictions on $L,c_{1},c_{2},Z$ and $k$ given in Theorems 5.10, 5.13 and 5.14, we show that $\ker(\mathcal{L}_{Z})=\{\mathbb{0}\}$ and that the Morse index of $\mathcal{L}_{Z}$ is equal to one. Therefore, as a consequence of the linear instability criterion in Theorem 3.2 we deduce that these single-lobe kink states are linearly unstable by the sine-Gordon flow on a tadpole graph (see our main instability result in Theorem 5.15 below).
@@
For completeness and for the convenience of the reader, in the Appendix we formulate some tools from the extension theory of symmetric operators by Krein and von Neumann which are suitable for our needs. In particular, we establish a Perron-Frobenius result which is used in the accurate Morse index estimate $n(\mathcal{L}_{Z})\leqq 1$ for $\mathcal{L}_{Z}$ in (1.10) and $(\phi_{1},\phi_{2})$ being a priori single-lobe kink state (see Theorem 5.4).
@@
Let $A$ be a closed densely defined symmetric operator in a Hilbert space $H$. The domain of $A$ is denoted by $D(A)$. The deficiency indices of $A$ are denoted by $n_{\pm}(A):=\dim\ker(A^{*}\mp iI)$, with $A^{*}$ denoting the adjoint operator of $A$. The number of negative eigenvalues counting multiplicities (or Morse index) of $A$ is denoted by $n(A)$. For any $\infty\leq a<b\leq\infty$, we denote by $L^{2}(a,b)$ the Hilbert space equipped with the inner product $(u,v)=\int_{a}^{b}u(x)\overline{v(x)}dx$. By $H^{m}(a,b)$ we denote the classical Sobolev spaces on $(a,b)\subseteq\mathbb{R}$ with the usual norm. We denote by $\mathcal{G}$ the tadpole graph parametrized by the edges $e_{1}=[-L,L]$, $e_{2}=[L,\infty)$, attached to a common vertex $\nu=L$. On the graph $\mathcal{G}$ we define the classical spaces $L^{p}(\mathcal{G})=L^{p}(-L,L)\oplus L^{p}(L,\infty)$, $p>1$, and
@@
$\langle\mathbf{u},\mathbf{v}\rangle=\int_{-L}^{L}u_{1}(x)\overline{v_{1}(x)}\,dx+\int_{L}^{\infty}u_{2}(x)\overline{v_{2}(x)}\,dx.$
Depending on the context we will use the following notations for different objects. By $\|\cdot\|$ we denote the norm in $L^{2}(\mathbb{R})$ or in $L^{2}(\mathcal{G})$. By $\|\cdot\|_{p}$ we denote the norm in $L^{p}(\mathbb{R})$ or in $L^{p}(\mathcal{G})$.